OpenStudy (anonymous):

Express the complex number in trigonometric form. 4i

hartnn (hartnn):

trigonometry form : r (cos theta +i sin theta) so, 3i = 0+3i now cos part is 0, and the angle which gives cos theta =0 is 90 degrees = 3 (cos 90 +i sin 90) thats it!

OpenStudy (anonymous):

Thanks for the help, hartnn! But where does the 4 go?

mathslover (mathslover):

He meant, trigonometry form : r (cos theta +i sin theta) so, 4i = 0+4i now cos part is 0, and the angle which gives cos theta =0 is 90 degrees = 4 (cos 90 +i sin 90) thats it!

hartnn (hartnn):

i solved for 3i you solve for 4i ^_^

OpenStudy (anonymous):

Ohhhh okay thanks guys! is there anyway to give medals to 2 people?

mathslover (mathslover):

No @kjuchiha , I didn't do anything here. Just give medal to Hartnn ... He deserves it , not I.

mathslover (mathslover):

btw, @hartnn instead of taking 90 degree or pi/2 , will it be better to consider principle value for theta?

OpenStudy (anonymous):

Alright man thanks guys

hartnn (hartnn):

give an example with principal value....

mathslover (mathslover):

Quest. : Express the complex number in trigonometric form. \(\cfrac{-1}{2} - \cfrac{i\sqrt{3}}{2}\) r cos theta = -1/2 r sin theta = \(- \cfrac{\sqrt{3}}{2}\) Therefore, r = 1 Also , tan theta = \(\sqrt{3}\) = \(\tan \cfrac{\pi}{3}\) Since, \((-1/2 , - \cfrac{\sqrt{3}}{2})\) lies in III quadrant. Therefore, Principle value of theta is \(-\pi + \cfrac{\pi}{3}\) Therefore, we get : \(\cos (-\cfrac{2\pi}{3} ) + i \sin ( - \cfrac{2\pi}{3} )\)

mathslover (mathslover):

Am I wrong somewhere, @hartnn ? Or is it right?

hartnn (hartnn):

i think its correct...in this case(4i) principle value = 90, right ?